irections. From your family members, form a group of five. (Just in case your



family is composed of less than five members, you can add one from your



neighbors). Get the weight in the kilogram of each member of the group. Draw



random samples of size 𝑛 = 2 from these weights.



1. List all possible samples and compute the mean of each sample.



2. Construct the sampling distribution of the sample means.



3. Find the mean of the population πœ‡.



4. Find the standard deviation of the population 𝜎.



5. Find the mean of the sampling distribution of the sample means πœ‡π‘‹Μ….



6. Find the standard deviation of the sampling distribution of the sample



means πœŽπ‘‹Μ….



7. Verify the Central Limit Theorem by:



a. Comparing πœ‡ and πœ‡π‘‹Μ….



b. Comparing 𝜎 and πœŽπ‘‹Μ….

1
Expert's answer
2022-03-17T18:10:22-0400

Solution

Population size "N=5"

Weights obtained "=52,61,68,78,84"

Sample size "n=2"


1. Possible samples

"=C_5^2= \binom{5}{2}=10" samples


"~~~~\def\arraystretch{1.5}\begin{array}{c:c:c}no
&Sample& Sample ~mean\\\hline1&52,61&56.5\\\hdashline2&52,68&60.0\\\hdashline3&52,78&65.0\\\hdashline4&52,84&68.0\\\hdashline5&61,68&64.5\\\hdashline6&61,78&69.5\\\hdashline7&61,84&72.5\\\hdashline8&68,78&73.0\\\hdashline9&68,84&76.0\\\hdashline10&78,84&81.0\\\hline\end{array}"


2. Sampling distribution of the sample means

"~~~~~\def\arraystretch{1.5}\begin{array}{c:c:c}mean&freq& prob\\\hline56.5&1&1/10\\\hdashline60.0&1&1/10\\\hdashline64.5&1&1/10\\\hdashline65.0&1&1/10\\\hdashline68.0&1&1/10\\\hdashline69.5&1&1/10\\\hdashline72.5&1&1/10\\\hdashline73.0&1&1/10\\\hdashline76.0&1&1/10\\\hdashline81.0&1&1/10\\\hline\end{array}"


3. Mean of the population.

"\mu=\dfrac{\sum X_i}{N}"


"\mu=\dfrac{52+61+68+78+84}{5}"


"\mu =68.6"


4. Standard deviation


"\sigma=\sqrt{\dfrac{\sum (X_i-\bar X)^2}{N}}"


"~~~\def\arraystretch{1.5}\begin{array}{c:c:c}X_i&X_i-\bar X&(X_i-\bar X)^2\\\hline56.5&-12.1&146.41\\\hdashline60.0&-8.6&73.96\\\hdashline64.5&-4.1&16.81\\\hdashline65.0&-3.6&12.96\\\hdashline68.0&-0.6&0.36\\\hdashline69.5&0.9&0.81\\\hdashline72.5&3.9&15.21\\\hdashline73.0&4.4&19.36\\\hdashline76.0&7.4&54.76\\\hdashline81.0&12.4&153.76\\\hline\sum&&494.40\\\hline\end{array}"


"\sigma=\sqrt{\dfrac{494.40}{5}}=9.944"


5. Mean of the sampling distribution

"\mu_x=\sum XP(X)"


"\mu_x=56.5(0.1)+60.0(0.1)+64.5(0.1)+65.0(0.1)+68.0(0.1)+69.5(0.1)+72.5(0.1)+73.0(0.1)+76.0(0.1)+81.0(0.1)"


"\mu_x=68.6"


6. Standard deviation of the sampling distribution of the sample means.

"\sigma_x=\dfrac{\sigma}{\sqrt N}=\dfrac{9.944}{\sqrt 5}"


"\sigma_x=4.45"


7. Verification

The central limit theory states that if you take samples of larger and larger size from any population, then the mean "\overline{x}" of the sample tends to get closer and closer to "\u03bc" and the standard deviation gets smaller.


(a) "\mu=\mu_{\overline x} ~~~(68.6=68.6)"

hence verified.


(b) "\sigma >\sigma_{\overline x}~~~(9.944>4.45)"

hence verified.